OpenStudy (anonymous):

how to solve for cosh(3i)?

OpenStudy (alekos):

coshx = (e^x + e^-x)/2 so just substitute x=3i

OpenStudy (anonymous):

oh thanks! :)

OpenStudy (anonymous):

can't solve it using a sci. calculator tho..

OpenStudy (alekos):

no you can't! need to convert e^3i to x+yi & e^-3i to u+vi

OpenStudy (danjs):

cosh(i*x) = cos(x)

OpenStudy (danjs):

using euler's formula

OpenStudy (anonymous):

convert? isn't 3i already in rectangular form? :)

OpenStudy (alekos):

no. e^3i is in polar form. but Dan has given it away

OpenStudy (danjs):

\[\cosh(i*x) = \frac{ 1 }{ 2 }[e ^{ix} + e ^{-ix}] = \cos(x)\]

OpenStudy (alekos):

e^3i = cos3 + isin3 and e^-3i = cos(-3) + isin(-3) so (e^3i + e^-3i)/2 = cos3

OpenStudy (danjs):

all u have to know is euler's formula

OpenStudy (alekos):

now you can use a calculator :)

OpenStudy (perl):

also you can use the fact that cos(-x) = cos x